# Quantum-Resistant RSA Modulus Decomposition via Adaptive Rényi Entropy Optimization

### v1

@@ VII Theoretical Extensions
This section explores advanced theoretical extensions of our framework, including hybrid constructions and information-theoretic security [16]. These extensions demonstrate the versatility and foundational significance of our entropy optimization approach.
VII-A Hybrid Cryptosystem Construction
We define a quantum-resistant hybrid scheme combining AREO-RSA with lattice-based cryptography [19]:
$\text{KeyGen}_{\text{hybrid}}(1^{k}):$
$\quad(pk_{\text{RSA}},sk_{\text{RSA}})\leftarrow\text{AREO-RSA.KeyGen}(1^{k})$
$\quad(pk_{\text{LWE}},sk_{\text{LWE}})\leftarrow\text{LWE.KeyGen}(1^{k})$
$\quad PK\leftarrow(pk_{\text{RSA}},pk_{\text{LWE}}),SK\leftarrow(sk_{\text{RSA}},sk_{\text{LWE}})$
Encryption encrypts message $m$ as:
$c=(\text{AREO-RSA.Enc}(K),\text{LWE.Enc}(m;K))$ (47)
where $K$ is random session key.
Theorem 10.
The hybrid scheme achieves CCA2 security under quantum random oracle model [20] if both components are IND-CPA secure.
This hybrid construction provides a practical migration path while maintaining strong security guarantees [19].
@@ VII-A Hybrid Cryptosystem Construction
By sequential composition theorem [23], adversary advantage decomposes as:
$\mathbf{Adv}_{\mathcal{A}}^{\text{CCA2}}$ $\leq\mathbf{Adv}_{\mathcal{B}}^{\text{AREO-IND}}+\mathbf{Adv}_{\mathcal{C}}^{\text{LWE-IND}}$
$+\text{negl}(k)$ (48)
where $\mathcal{B},\mathcal{C}$ are reduction adversaries. Since both components reduce to hard problems (IFP and LWE) [12], the hybrid achieves CCA2 security [23]. ∎
@@ VII-B Quantum Information-Theoretic Security
We extend our framework to information-theoretic security using quantum information theory [16], providing unconditional protection against quantum adversaries. The following definition provides a fundamental security metric independent of computational assumptions [24].
@@
where $P_{e}$ is measurement error probability. Set $P_{e}<\epsilon/\log k$ and apply Holevo bound [15]:
@@
This corollary directly connects our control parameter $\gamma$ to information-theoretic security, enabling parameter selection for unconditional security [16].
@@ VIII Conclusion
We summarize our core contributions, highlight the theoretical advances, and outline promising research directions. We have established a rigorous theoretical foundation for enhancing RSA’s quantum resistance through adaptive Rényi entropy optimization. Key contributions include:
- Formal proof of $\Omega(2^{k/3})$ quantum attack complexity via eigenvalue degeneracy arguments [14] - Construction of entropy-constrained prime generation framework with polynomial-time implementation - Security reduction to Ring-LWE [12] with explicit lattice embedding - Quantum information-theoretic security framework via Fano bound [16] - Classical security preservation against Wiener [4], Boneh-Durfee [5], and Coppersmith attacks [6]
Theoretical Breakthroughs: 1. Resolved prime existence problem for $\gamma<2^{-k/6}$ via Bombieri-Vinogradov theorem [13] 2. Constructed explicit lattice embedding $\phi:\mathbb{Z}_{N}^{*}\hookrightarrow R_{q}^{\vee}$ 3. Established quantum Fano bound for unconditional security [16] 4. Unified classical and quantum security through entropy constraints [14]
Compatibility Note: While our prime proximity constraint $|p-q|<\gamma\sqrt{pq}$ differs from PKCS #1’s recommendation $|p-q|>2^{k/2-100}$ [2], this modification is essential for quantum resistance. Standard compliance can be maintained through protocol-level adjustments preserving the entropy optimization benefits.
The theoretical framework opens several research directions:
@@
    Optimizing entropy bounds via zeta function distributions
@@
    Exploring connections to Riemann hypothesis in prime gap analysis [13]
@@
    Extending the framework to multi-prime RSA variants [7]
@@
    Investigating quantum query complexity limits [14]
@@
    Developing formal verification methods for entropy-constrained primes
Acknowledgments
This research received no external funding. The authors thank the anonymous reviewers for their valuable insights and constructive feedback, which significantly improved the theoretical rigor of this work. Special thanks to Prof. Jianfeng Ma for insightful discussions on quantum information theory.
Biography Section
Ruopengyu Xu is an independent researcher focusing on post-quantum cryptography and quantum information theory. He holds a bachelor’s degree. His research interests lie at the intersection of cryptography, number theory, and quantum information.
Chenglian Liu is a professor at the School of Electrical and Computer Engineering of Nanfang College Guangzhou. He holds a PhD from Royal Holloway, University of London. His research interests include cryptography, number theory and their integrated applications.

### v2

@@ VII Theoretical Extensions
This section explores theoretical extensions of our framework.
VII-A Standard-Compatible Construction
To maintain compatibility with existing standards requiring $|p-q|>N^{1/4}$, we propose a layered approach:
Theorem 10 (Compatible Prime Construction).
There exist primes $p^{\prime},q^{\prime}$ satisfying both:
  1. 1.
    $|p^{\prime}-q^{\prime}|>2^{k/2-100}$ (PKCS#1 requirement [2])
  2. 2.
    Inner primes $p,q$ with $|p-q|<\gamma\sqrt{pq}$ (quantum resistance)
via the construction:
$p^{\prime}$ $=2^{100}\cdot K+p$ (52)
$q^{\prime}$ $=2^{100}\cdot(K+\delta)+q$ (53)
where $K$ is a $\lfloor k/2\rfloor$-bit integer, $\delta\geq 1$, and $p,q$ are $\lfloor k/2\rfloor$-bit primes satisfying proximity.
@@ VII-A Standard-Compatible Construction
The outer difference satisfies:
$|p^{\prime}-q^{\prime}|\geq 2^{100}|\delta|-|p-q|>2^{100}-2^{k/2}\gg 2^{k/2-100}$ (54)
for $k>2048$. The modulus $N^{\prime}=p^{\prime}q^{\prime}$ preserves the entropy properties of the inner primes because:
$a^{x}\mod N^{\prime}\equiv a^{x}\mod p\mod p$ (55)
so the period-finding operator has similar spectral properties as for $N=pq$.
The Fermat factorization method [28] requires:
$|p^{\prime}-q^{\prime}|<N^{1/4}$ (56)
which is not satisfied by our construction since $|p^{\prime}-q^{\prime}|>2^{100}>2^{k/2-100}>N^{1/4}$ for $k>2048$. ∎
@@ VII-B Quantum Information-Theoretic Security
We extend our framework to information-theoretic security.
@@
Set $P_{e}<\epsilon/\log k$ and apply Holevo bound [13]:
@@
VII-C Multi-Prime RSA Extension
We extend the framework to multi-prime RSA [29] with modulus $N=p_{1}p_{2}\cdots p_{m}$:
Theorem 12 (Multi-Prime Entropy Bound).
For $m$-prime RSA with primes satisfying $\max_{i,j}|p_{i}-p_{j}|<\gamma\sqrt{N}$, the Rényi entropy satisfies:
$\mathscr{H}_{2}(\rho_{N})\leq\log m+2\log\left(1+\frac{\gamma}{2}\right)+\mathcal{O}(k^{-1/2})$ (62)
Proof.
The quantum state for modulus $N$ has at least $m$ significant eigenvalues. The purity is bounded by:
$\operatorname{Tr}(\rho^{2})\leq\lambda_{\max}^{2}+\sum_{i=2}^{m}\lambda_{i}^{2}+\sum_{j>m}\lambda_{j}^{2}$ (63)
where $\lambda_{\max}\leq\frac{1}{m}+\mathcal{O}(\gamma)$. The result follows from eigenvalue concentration. ∎
Algorithm 2 Multi-Prime AREO-RSA Key Generation
0: Security parameter $k$, prime count $m$, proximity $\gamma$
0: Public key $(N,e)$, private key $(d,p_{1},\dots,p_{m})$
1: Generate $m$ primes $p_{1},\dots,p_{m}$ satisfying:
  • $|p_{i}-p_{j}|<\gamma\sqrt{p_{i}p_{j}}$ for all $i,j$
  • $p_{i}\equiv a_{i}\mod M$ for distinct residues [8]
2: Compute $N=\prod_{i=1}^{m}p_{i}$, $\phi(N)=\prod_{i=1}^{m}(p_{i}-1)$
3: Select $e$ coprime to $\phi(N)$, compute $d=e^{-1}\mod\phi(N)$
4:return  $(N,e),(d,p_{1},\dots,p_{m})$
This extension demonstrates the flexibility of our entropy optimization framework to various RSA variants.
VII-D Relaxed Entropy Constraints
For practical implementation, we can relax the entropy constraints while maintaining security:
Lemma 3 (Polynomial Entropy Bound).
When $\gamma>\frac{\log k}{\sqrt{k}}$, the existence probability satisfies:
$\mathbb{P}(\exists p,q\text{ satisfying constraints})>k^{-c}$ (64)
for some constant $c>0$.
Proof.
Apply the prime number theorem with error term $O(e^{-c\sqrt{\log N}})$. For $\gamma=\Omega(\log k/\sqrt{k})$, the density in Theorem 3 becomes:
$\pi_{2}(N,\gamma)\gg\frac{N}{\log^{2}N}\cdot\frac{\log k}{\sqrt{k}}=\frac{N\log k}{k^{1/2}\log^{2}N}$ (65)
which is $\Omega(N^{1-\epsilon})$ for any $\epsilon>0$ when $k$ is large. Thus the existence probability is at least $k^{-c}$ for sufficiently large $k$.
The Heath-Brown prime gap estimate [30] provides:
$\sum_{n\leq x}\Lambda(n)\Lambda(n+h)=\mathfrak{S}(h)x+O(x^{5/6})$ (66)
for $h\leq x^{5/6}$, supporting our density claim for polynomial $\gamma$. ∎
@@ VIII Conclusion
We have explored a theoretical approach to enhance RSA’s quantum resistance through Rényi entropy optimization. Key aspects include:
- Established connection between prime proximity and quantum complexity via Maynard’s theorem [18] - Constructive existence proofs for entropy-constrained primes [22] - Security reduction to ideal lattice problems [16, 27] - Compatibility-preserving constructions for existing standards [28] - Extension to multi-prime RSA variants [29]
Theoretical contributions: 1. Prime gap analysis via Maynard’s theorem [18] 2. Ideal lattice embedding for SVP reduction [16] 3. Quantum information-theoretic bounds [14] 4. Standard-compatible layered construction [28] 5. Multi-prime RSA extension with entropy constraints [29]
Future research may explore:
@@
    Tight bounds for quantum complexity enhancement [19]
@@
    Concrete security parameters for practical deployment [17]
@@
    Extension to ring-based cryptosystems [16]
@@
    Quantum query complexity limits [19]
@@
    Verification protocols for entropy constraints
  6. 6.
    Implementation challenges and optimizations [31]
This work provides a theoretical foundation for enhancing RSA’s quantum resistance while preserving its algebraic structure, though practical realization requires further investigation.